OpenStudy (anonymous):

What is the solution of the following system? {-2x-y=1 {-4x-2y=-1

OpenStudy (anonymous):

you can make it 2x+y=-1 and 4x+2y=1

OpenStudy (anonymous):

then if you add them, it makes 6x+3y=0

OpenStudy (anonymous):

divide both sides by 3 to make it 2x+y=0

OpenStudy (anonymous):

wait scratch that